# Tree metrics and log-concavity for matroids

### v1

@@ 3.2. Lorentzian property of the homogeneous generating polynomial $Z_{q,\nu}$
The following lemma is a variation of known relations between tree metrics, ultrametrics, and valuated matroids. See, for instance, [13, Section 4.3].
@@
By the local exchange property for $\text{M}^{\natural}$-concave functions [16, Theorem 6.4], for any distinct elements $i,j,k$ in $E$, the maximum among the three numbers
@@
The “if” direction is [5, Proposition 5.5]. We show that $Z_{q,\nu}$ is Lorentzian for positive $q\leq 1$ when $\nu$ is $\text{M}^{\natural}$-concave.
@@
This is a special case of [4, Lemma 3.27], because restrictions of M-concave functions to coordinate half-spaces are M-concave [16, Section 6.4]. Since $\mathrm{L}^{d}_{n}$ is closed, we may assume that $0<q<1$, and for any such $q$, we have
@@
where $\nu/i$ is the $\text{M}^{\natural}$-concave function on $2^{E\setminus i}$ defined by $\nu/i(S)=\nu(S\cup i)$. For general discussion about the contraction $\nu/i$ of an $\text{M}^{\natural}$-concave function $\nu$, see [16, Section 6.4]. Thus, modulo induction on $n$, we only need to consider the Lorentzian property of the quadratic form
@@
Identifying the variables $x_{i}$ with each other in $Z_{q,\nu}(\mathsf{x},y)$, we get a bivariate polynomial with coefficients $I_{q,\nu;k}$. Since $Z_{q,\nu}$ is Lorentzian by Theorem 1.5, the specialization is Lorentzian as well [4, Theorem 2.10], and hence the sequence $I_{q,\nu;k}$ is ultra log-concave in $k$ for positive $q\leq 1$ [4, Example 2.3]. ∎
@@ 3.3. Polynomial log concavity
The polynomial $\partial_{1}^{i}f$ is Lorentzian [4, Corollary 2.11], and hence its specialization $i!f_{i}$ is Lorentzian as well [4, Theorem 2.10]. ∎
@@
Fix a positive real parameter $q\leq 1$, and consider two copies of homogeneous generating polynomial of an $\text{M}^{\natural}$-concave function $\nu$, say $Z_{q,\nu}(\mathsf{x},y)$ and $Z_{q,\nu}(\mathsf{x},z)$, where we use distinct homogenizing variables $y$ and $z$ for the same set of $n$ variables $\mathsf{x}$. Since $Z_{q,\nu}(\mathsf{x},y)$ and $Z_{q,\nu}(\mathsf{x},z)$ are Lorentzian by Theorem 1.5, their product is Lorentzian too [4, Corollary 2.32]. Writing $n$ for the cardinality of $E$ as before, we have
@@
is Lorentzian. Therefore, the sequence of normalized coefficients $(n-j)!\ (n-k)!\ N^{S}_{q,\nu}(j,k)$ is log-concave and has no internal zeros [4, Example 2.3]. It follows that, for all $i\leq j\leq k\leq l$ for with $i+l=j+k=\left|S\right|$, we have

### v2

@@ 3.2. Lorentzian property of the homogeneous generating polynomial $Z_{q,\nu}$
The following lemma is a variation of known relations between tree metrics, ultrametrics, and valuated matroids. See, for instance, [MS15, Section 4.3].
@@
By the local exchange property for $\text{M}^{\natural}$-concave functions [Mur03, Theorem 6.4], for any distinct elements $i,j,k$ in $E$, the maximum among the three numbers
@@
The “if” direction is [EH20, Proposition 5.5]. We show that $Z_{q,\nu}$ is Lorentzian for positive $q\leq 1$ when $\nu$ is $\text{M}^{\natural}$-concave.
@@
This is a special case of [BH20, Lemma 3.27], because restrictions of M-concave functions to coordinate half-spaces are M-concave [Mur03, Section 6.4]. Since $\mathrm{L}^{d}_{n}$ is closed, we may assume that $0<q<1$, and for any such $q$, we have
@@
where $\nu/i$ is the $\text{M}^{\natural}$-concave function on $2^{E\setminus i}$ defined by $\nu/i(S)=\nu(S\cup i)$. For general discussion about the contraction $\nu/i$ of an $\text{M}^{\natural}$-concave function $\nu$, see [Mur03, Section 6.4]. Thus, modulo induction on $n$, we only need to consider the Lorentzian property of the quadratic form
@@
Identifying the variables $x_{i}$ with each other in $Z_{q,\nu}(\mathsf{x},y)$, we get a bivariate polynomial with coefficients $I_{q,\nu;k}$. Since $Z_{q,\nu}$ is Lorentzian by Theorem 1.5, the specialization is Lorentzian as well [BH20, Theorem 2.10], and hence the sequence $I_{q,\nu;k}$ is ultra log-concave in $k$ for positive $q\leq 1$ [BH20, Example 2.3]. ∎
@@ 3.3. Polynomial log concavity
The polynomial $\partial_{1}^{i}f$ is Lorentzian [BH20, Corollary 2.11], and hence its specialization $i!f_{i}$ is Lorentzian as well [BH20, Theorem 2.10]. ∎
@@
Fix a positive real parameter $q\leq 1$, and consider two copies of homogeneous generating polynomial of an $\text{M}^{\natural}$-concave function $\nu$, say $Z_{q,\nu}(\mathsf{x},y)$ and $Z_{q,\nu}(\mathsf{x},z)$, where we use distinct homogenizing variables $y$ and $z$ for the same set of $n$ variables $\mathsf{x}$. Since $Z_{q,\nu}(\mathsf{x},y)$ and $Z_{q,\nu}(\mathsf{x},z)$ are Lorentzian by Theorem 1.5, their product is Lorentzian too [BH20, Corollary 2.32]. Writing $n$ for the cardinality of $E$ as before, we have
@@
is Lorentzian. Therefore, the sequence of normalized coefficients $(n-j)!\ (n-k)!\ N^{S}_{q,\nu}(j,k)$ is log-concave and has no internal zeros [BH20, Example 2.3]. It follows that, for all $i\leq j\leq k\leq l$ for with $i+l=j+k=\left|S\right|$, we have